To burn the fats she accumulated during the enhanced quarantine time, Inday walks south at a speed of 3 m/s for 10 minutes. She then turns around and walks north a distance 1000 m in 5 minutes. What is her total distance covered?


a. 2,800 m

b. 11,800 m

c. 300,000 m

d. 2,030 m


1
Expert's answer
2022-02-10T09:18:09-0500

Given:

"v_1=\rm 3\: m/s"

"t_1=\rm 10\: min=600\: s"

"t_2=\rm 5\: min=300\: s"

"d_2=1000\:\rm m"


The total displacement

"d=d_1+d_2=v_1t_1+d_2\\
=\rm 3\: m/s*600\: s+1000\: m=2800\: m"

Answer: A